Compute the z-statistic for each given claim (p), the observed proportion (p^), and the given



sample size (𝑛).




𝑝 = 0.66, p^= 0.61, 𝑛 = 40

1
Expert's answer
2022-05-30T07:44:38-0400
"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.61-0.66}{\sqrt{\dfrac{0.66(1-0.66)}{40}}}=-4.5394"